# Towards Federated Learning with on-device Training and Communication in 8-bit Floating Point

### v1

@@ Abstract.
Recent work has shown that 8-bit floating point (FP8) can be used for efficiently training neural networks with reduced computational overhead compared to training in FP32/FP16. In this work, we investigate the use of FP8 training in a federated learning context. This brings not only the usual benefits of FP8 which are desirable for on-device training at the edge, but also reduces client-server communication costs due to significant weight compression. We present a novel method for combining FP8 client training while maintaining a global FP32 server model and provide convergence analysis. Experiments with various machine learning models and datasets show that our method consistently yields communication reductions of at least 2.9x across a variety of tasks and models compared to an FP32 baseline.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 25, 2024; Barcelona, Spainisbn: 978-x-xxxx-XXXX-X/YY/MM
1. Introduction
@@
Lots of data is generated daily on personal smartphones and other devices at the edge. This data is very valuable for training machine learning models to provide services such as better voice recognition or text completion. However, the local data carries sensitive personal information which needs to be protected for privacy reasons. Furthermore, communication of local data between billions of devices and data centers is expected to occupy lots of network bandwidth and transmission is costly in terms of power consumption, which is a primary concern for edge devices running on batteries.
@@
Despite these constraints, it is still possible to train a model without having to transmit this local data using federated learning (McMahan et al., 2017). In federated learning, each local device performs training locally with their local data and update their local models. When it comes to communication, the central server receives local models from a subset of devices. The central server then aggregates these local models and transmits a new global model back to those devices for a model update. In this way, no local data is ever exposed during communication and the global model can learn from local data as communication goes on.
@@
Since its inception, new techniques around federated learning have been proposed to reduce communication cost. The local models, albeit smaller than the local data, are still expensive to transmit via wireless communication and will be taxing on local devices’ battery life if performed very frequently. One method to reduce communication cost is to quantize the models before the communication occurs, and several works have shown that this can be done without significant loss in model accuracy (Gupta et al., 2023; Hönig et al., 2022; Ni et al., 2024). Nevertheless, standard quantization of the model weights will introduce a bias term that slows down convergence of the training process. In order to alleviate this problem, the use of stochastic rounding has been proposed (Zheng et al., 2020; Bouzinis et al., 2023). Hence, when aggregating the client models at the server, the stochastic rounding errors tend to zero as the number of clients grows large. This method has also been shown to improve the learning process from a privacy perspective (Youn et al., 2023). Furthermore, He et al. (He et al., 2020) showed that stochastic rounding in conjunction with non-linear quantization can reduce the number of communication rounds to reach convergence even more.
Figure 1. Overview of federated learning with local FP8 on-device training and weight quantization in both uplink and downlink communication.
@@
While integer representations, such as INT8, have been widely adopted for neural network quantization for efficient inference, the use of FP8 remains relatively new in comparison and has been more focused on model training. A few recent works (Wang et al., 2018; Sun et al., 2019; Peng et al., 2023) have proposed centralized neural network training in FP8 with promising results. However, for some networks, the FP8 precision was found to be insufficient for retaining accuracy in certain operations. Notably, the backwards pass through the network typically requires higher dynamic range than the forward pass. To this end, Micikevicius et al. (Micikevicius et al., 2022) proposed a binary interchange FP8 format that uses both E4M3 (4-bit exponent and 3-bit mantissa) and E5M2 (5-bit exponent and 2-bit mantissa) representations, which allows for minimal accuracy drops compared to FP16 across a wide range of network architectures. Concurrent work by Kuzmin et al. (Kuzmin et al., 2022) proposed a similar solution, where the exponent bias is flexible and updated for each tensor during training, which allows for maintaining different dynamic ranges in different parts of the network.
@@
Being a very efficient training datatype, FP8 is a good candidate for on-device training at the edge and its industrial support points to wide-spread real-world applications. A future scenario where edge devices can perform efficient on-device training with native hardware support introduces a new class of federated learning problems. It also increases device heterogeneity in a federated learning setting, where participating devices and servers may have different levels of hardware support for FP8.
@@
In this work, we introduce an implementation of federated learning with on-device FP8 training simulated by quantization-aware training (QAT), which learns from quantized models effectively while being efficient in its communication and computing cost. A high-level overview is shown in Figure 1. We present theoretical convergence properties as well as thorough experiments on a variety of models and data sets. Finally, we also present a novel server-side weight aggregation method that improves the server model accuracy compared to standard federated averaging.
2. Method
@@
Preliminaries. Consider the federated learning problem, where $K$ clients update their local models by training on disjoint local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$. Each client minimizes their own local objectives $F_{k}(\mathbf{w},Q,\alpha,\beta)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{k}}[l(\mathbf{w};\mathbf{x},y,Q,\alpha,\beta)]$, where $Q$ is a quantization operator and $l$ is the loss function. Furthermore, $\alpha$ and $\beta$ are the per-tensor clipping values used for weights and activations respectively. Henceforth, we denote quantized weights based on range $\alpha$ as $Q(\mathbf{w};\alpha)$. In practice, different clipping values are used for different layers of the network, but we omit this in our notation for readability. The overall problem can be expressed as
(1) $\min_{\mathbf{w}}F(\mathbf{w},Q,\alpha,\beta),\quad F(\mathbf{w},Q,\alpha,\beta)=\sum_{k=1}^{K}\frac{n_{k}}{n}F_{k}(\mathbf{w},Q,\alpha,\beta),$
@@
On-Device Quantization-Aware Training. Depending on the hardware support, on-device local training can be performed in native FP8 or using quantization-aware training (QAT), or a mix of the two. Native FP8 training is supported by the latest AI hardware in data centers such as Nvidia’s H100/H200 series of GPUs. There is significant industry support behind FP8, it is only a matter of time for FP8 hardware support to arrive on edge devices. For research purposes, we here resort to QAT, and follow the FP8 QAT method described in (Kuzmin et al., 2022), using per-tensor quantization for both model weights and activations, with one signed bit, and $m=3$ and $e=4$ bits for the mantissa and exponent respectively, as well as a flexible exponent bias. QAT with both weights and activations quantization is a good way of simulating native FP8 training on supported hardware with low precision arithmetic. In our setting, we are not simulating the effect of gradient quantization which happens on FP8-enabled hardware. However, previous work (Kuzmin et al., 2022) has shown that it is a good approximation to ignore its effect.
@@
Let $\mathbf{x}=(x_{1},...,x_{d})^{T}\in\mathbb{R}^{d}$ denote an FP32 input tensor and $Q_{\text{det}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ the FP8 deterministic quantization operator with a clipping parameter $\alpha$, whose element-wise outputs are given by $Q_{\text{det}}(x_{i};\alpha)=s_{i}\left\lfloor\frac{x_{i}}{s_{i}}\right\rceil$. Here, $s_{i}$ is the scale that is computed as
@@
(2) $\log_{2}s_{i}=\begin{cases}\left\lfloor\log_{2}|x_{i}|+b\right\rfloor-b-m,\quad&\left\lfloor\log_{2}|x_{i}|+b\right\rfloor>1\\ 1-b-m,\quad&\text{otherwise,}\end{cases}$
@@
where the exponent bias depends on the clipping value $\alpha$ as $b=2^{e}-\log_{2}\alpha+\log_{2}(2-2^{-m})-1$. At training time, $\alpha$ is first initialized using the maximum absolute value of each weight range, and then treated as a learnable parameter that is updated during training. Furthermore, the gradients of the non-differentiable rounding operators are computed using the straight-through estimator (Bengio et al., 2013), i.e. $\frac{\partial\left\lfloor x_{i}\right\rceil}{\partial x_{i}}=1$, with a key exception being $\left\lfloor\log_{2}|x_{i}|+b\right\rfloor$, which is treated as a constant following the approach in Kuzmin et al. (2022). Activations are quantized using the same procedure, but with a separate clipping value $\beta$.
@@
Unbiased Quantized Communication. When applying FP8 QAT to a federated leaning scenario, an important aspect is the ability to reduce communication overhead by transferring weights between clients and the server using only 8 bits per scalar value. On client devices with hardware for FP8 mixed-precision training support, a copy of high-precision master weights (Vishwanath, 2019) is present as in our QAT setup. Therefore, at the end of each communication round, the participating clients need to perform a hard reset of their master weights to the de-quantized FP8 values on a quantization grid. This approach allows for cost reduction in both the uplink and downlink communication.
@@
(3) $Q_{\text{rand}}(x_{i};\alpha)=s_{i}\begin{cases}\left\lceil\frac{x_{i}}{s_{i}}\right\rceil&p\leq\frac{x_{i}}{s_{i}}-\left\lfloor\frac{x_{i}}{s_{i}}\right\rfloor\\ \left\lfloor\frac{x_{i}}{s_{i}}\right\rfloor&\text{otherwise,}\end{cases}$
@@
where $p$ is a Bernoulli random variable that takes the values 0 and 1 with equal probability. It is straightforward to verify that this randomized quantization is unbiased from a statistics point of view while the deterministic quantization introduced earlier is biased.
@@
The quantized weights are then aggregated at the server using a federated average as $\mathbf{w}_{t+1}\leftarrow\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1})$, where $m_{t}=\sum_{{k^{\prime}}\in\mathcal{P}_{t}}n_{k^{\prime}}$. The clipping values are also aggregated, but without quantization, their contribution to the communication overhead is small relative to the weights. The aggregated weights are then quantized again to FP8 and transmitted to the next set of active clients with a new set of quantization parameters.

### v2

@@ Abstract
Recent work has shown that 8-bit floating point (FP8) can be used for efficiently training neural networks with reduced computational cost compared to training in FP32/FP16. In this work, we investigate the use of FP8 training in a federated learning context. This approach brings not only the usual benefits of FP8 which are desirable for on-device training at the edge, but also reduces client-server communication costs due to significant weight compression. We present a novel method for combining FP8 client training while maintaining a global FP32 server model and provide convergence analysis. Experiments with various machine learning models and datasets show that our method consistently yields communication reductions of at least 2.9x across a variety of tasks and models compared to an FP32 baseline to achieve the same trained model accuracy.
Keywords: federated learning, quantization, FP8
1 Introduction
@@
A large amount of data is generated daily on personal smartphones and other devices at the edge. This data is very valuable for training machine learning models to provide services such as better voice recognition leroy2019federated or text completion hard2018federated . However, the local data often carries sensitive personal information which needs to be protected for privacy reasons. Furthermore, communication of local data between billions of devices and data centers is expected to occupy lots of network bandwidth and transmission is costly in terms of power consumption, which is a primary concern for edge devices running on batteries.
@@
Despite these constraints, it is still possible to train a model without having to transmit this local data using federated learning mcmahan2017communication . In federated learning, each local device performs training locally with their local data and update their local models. When it comes to communication, the central server receives local models from a subset of devices. The central server then aggregates these local models and transmits a new global model back to those devices for a model update. In this way, no local data is ever exposed during communication and the global model can learn from local data as communication goes on.
@@
Since its inception, new techniques around federated learning have been proposed to reduce communication cost. The local models, albeit smaller than the local data, are still expensive to transmit via wireless communication and will be taxing on local devices’ battery life if performed very frequently. One method to reduce communication cost is to quantize the models before the communication occurs, and several works have shown that this can be done with limited reduction of model accuracy gupta2023quantization ; honig2022dadaquant ; ni2024fedaqt .
Figure 1: Overview of the proposed federated learning with local FP8 on-device training and weight quantization in both uplink and downlink communication. For time step $t$, the server sends quantized FP8 global model weights $Q_{\text{FP8}}(\mathbf{w}_{t})$ and range parameters $\alpha_{t}$, $\beta_{t}$ for the weights and activations respectively. Each client $k$ that participates in the training round then performs FP8 local training and sends back updated weight and range parameters. At the server, new global weights are obtained using global aggregation and an additional optimization step.
@@
Being a very efficient training datatype, FP8 is a good candidate for on-device training at the edge and the wide industrial support micikevicius2022fp8 behind it points to wide-spread real-world applications. A future scenario where edge devices can perform efficient on-device training with native hardware support introduces a new class of federated learning problems. It also increases device heterogeneity in a federated learning setting, where participating devices and servers may have different levels of hardware support for FP8.
@@
In this work, we introduce an implementation of federated learning with on-device FP8 training, which learns from quantized models effectively while being efficient in its communication and computing cost. A high-level overview of this method is shown in Figure 1. We summarize our main contributions as follows:
  1. 1.
    A novel method for combining local FP8 client training with an FP32 server model. The local FP8 training is simulated by quantization-aware training (QAT) and all communication between the server and the clients is done in FP8. Furthermore, we provide an additional optimization procedure for weight aggregation on the server that alleviates potential performance drops caused by quantization, without affecting communication costs.
  2. 2.
    We provide convergence analysis and motivation for the use of stochastic quantization for communication, but deterministic quantization for QAT. These design choices are further supported by experimental ablation studies.
  3. 3.
    We proved experiments on image and audio classification benchmarks, showing minimal loss in performance while obtaining large reductions in communication costs for both convolutional and transformer-based models.
2 Related Work
2.1 Federated Learning with Quantized Communication
Quantizing model weights is an effective way of reducing the communication cost of federated learning. For example, quantizing weights from 32 to 8 bits immediately reduces the number of communicated bits by 75% per training round. However, in practice the same functional performance is often not reached by the quantized model, because quantization can cause slower convergence of the training process. As a consequence, the communication reduction obtained to yield a particular performance threshold is often far from the ideal.
An important reason for the slower convergence obtained with quantized communication is that quantization of the model weights will introduce a bias term, which makes the server model a biased estimate of the average client model. In order to alleviate this problem, the use of stochastic rounding has been proposed zheng2020design ; bouzinis2023wireless . Hence, when aggregating the client models at the server, the stochastic rounding errors tend to zero as the number of clients grows large. This method has also been shown to improve the learning process from a privacy perspective youn2023randomized . Similar work, he2020cossgd , has shown that stochastic rounding in conjunction with non-linear quantization can reduce the number of communication rounds to reach convergence even more. Nevertheless, there is a limited amount of research on how to effectively combine quantized communication with low-precision client training.
yoon2022bitwidth investigated a similar setup as ours, i.e. low-precision local training with quantized communication, but only for uniform quanization schemes, such as INT4 and INT8. This leads to sub-optimal performance compared to full-precision training, since many neural networks are known to be sensitive to integer quantization, especially gradients during training are thought to require a bigger dynamic range kuzmin2022fp8 . For this reason, the industry is coalescing around FP8 for efficient training hardware micikevicius2022fp8 . This motivates an investigation of local training in low-precision floating point, which provides better training dynamics at the same communication cost.
2.2 FP8 Quantization for Neural Networks
While integer representations, have been widely adopted for neural network quantization for efficient inference, the use of FP8 remains relatively new in comparison and has been more focused on model training. A few recent works wang2018training ; sun2019hybrid ; peng2023fp8 have proposed centralized neural network training in FP8 with promising results. However, for some networks, a single FP8 representation was found to be insufficient for retaining accuracy in certain operations. Notably, the backwards pass through the network typically requires higher dynamic range than the forward pass. To this end, a binary interchange FP8 format that uses both E4M3 (4-bit exponent and 3-bit mantissa) and E5M2 (5-bit exponent and 2-bit mantissa) representations has been proposed, which allows for minimal accuracy drops compared to FP16 across a wide range of network architectures micikevicius2022fp8 . Concurrent work, kuzmin2022fp8 , proposed a similar solution, where the exponent bias is flexible and updated for each tensor during training, which allows for maintaining different dynamic ranges in different parts of the network.
An important factor in neural network quantization that is often overlooked, is that not all model architectures are equally sensitive to quantization, since the distribution of weights and activations are often architecture-dependent. As a consequence, a quantization scheme that works well for e.g. convolutional networks might not be suitable for attention-based models. shen2024efficient provided a thorough investigation of different quantization schemes for models on a wide variety of applications and found that FP8 quantization, with a combination of E4M3 and E3M4 in particular, was least detrimental to model performance. Similar results were found by nikolic2024schrodinger by dynamically adjusting the number of exponent and mantissa bits during training, which lead to large reductions in both memory footprint and energy consumption. This makes FP8 training particularly interesting in a federated training setup, where the available hardware resources are often limited.
3 Method
3.1 Preliminaries
@@
Consider the federated learning problem, where $K$ clients update their local models by training on disjoint local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$. Each client minimizes their own local objective functions $F_{k}(\mathbf{w},Q,\alpha,\beta)=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{k}}[l(\mathbf{w};\mathbf{x},y,Q,\alpha,\beta)]$, where $\mathbf{w}$ are the model parameters, $Q$ is a quantization operator and $l$ is the loss function. Furthermore, $\alpha$ and $\beta$ are the per-tensor clipping values used for maintaining the dynamic range when quantizing weights and activations respectively. Henceforth, we denote quantized weights based on range $\alpha$ as $Q(\mathbf{w};\alpha)$. In practice, each layer of the network has its own unique clipping parameters for both the weights and activations, but we omit this in our notation for readability.
We consider a modification of the standard federated averaging (FedAvg) mcmahan2017communication with quantized weights, where the objective is to find
$\min_{\mathbf{w}}$ $F(\mathbf{w},Q,\alpha,\beta),$ (1)
$F(\mathbf{w},Q,\alpha,\beta)$ $=\sum_{k=1}^{K}\frac{n_{k}}{n}F_{k}(\mathbf{w},Q,\alpha,\beta),$
@@
3.2 Floating Point Representation
A floating point number $x$ with $e$ exponent bits and $m$ mantissa bits can be written as
$x=(-1)^{s}2^{p-b}\left(1+\frac{d_{1}}{2}+\frac{d_{2}}{2^{2}}+\ldots+\frac{d_{m}}{2^{m}}\right),$ (2)
where $s\in\{0,1\}$ is the sign bit, $d_{i}\in\{0,1\}$ is the mantissa, $p\in\{0,1,\ldots,2^{e-1}\}$ is the exponent and $b$ is the exponent bias. In addition, we assume that $p=0$ is reserved for subnormal numbers, which allows an exact representation of 0 and other special values.
Compared to integer quantization, the quantization grid for floating point numbers is not uniform. Increasing the number of exponent bits results in a higher dynamic range, whereas increasing the number of mantissa bits increases the precision. Therefore, give a fixed budget on the number of bits to allocate, there is a trade-off to be made between the two.
3.3 On-Device Quantization-Aware Training
@@
Depending on the hardware support, on-device local training can be performed in native FP8 or using quantization-aware training (QAT), or a mix of the two. Native FP8 training is supported by the latest AI hardware in data centers such as Nvidia’s H100/H200 series of GPUs. There is significant industry support behind FP8 and it is only a matter of time for FP8 hardware support to arrive on edge devices.
For research purposes, we here resort to QAT, and follow the FP8 QAT method described in kuzmin2022fp8 , using per-tensor quantization for both model weights and activations, with one signed bit, and $m=3$ and $e=4$ bits for the mantissa and exponent respectively, as well as a flexible exponent bias that depends on learnable clipping parameters. QAT with both weights and activations quantization is a good way of simulating native FP8 training on supported hardware with low precision arithmetic. In our setting, we are not simulating the effect of gradient quantization which happens on FP8-enabled hardware. However, previous work, kuzmin2022fp8 , has shown that it is a good approximation to ignore its effect.
@@
Let $\mathbf{x}=(x_{1},...,x_{d})^{T}\in\mathbb{R}^{d}$ denote an FP32 input tensor and $Q_{\text{det}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ the FP8 deterministic quantization operator with a clipping parameter $\alpha$, whose element-wise outputs are given by
@@
$Q_{\text{det}}(x_{i};\alpha)=\begin{cases}-\alpha,\quad x_{i}\leq-\alpha,\\ s_{i}\left\lfloor\frac{x_{i}}{s_{i}}\right\rceil,\\ \alpha,\quad x_{i}\geq\alpha,\\ \end{cases}$ (3)
[…]
$p_{i}=\begin{cases}\left\lfloor\log_{2}|x_{i}|+b\right\rfloor-b-m,&\left\lfloor\log_{2}|x_{i}|{+}b\right\rfloor>1\\ 1-b-m,&\text{otherwise,}\end{cases}$ (4)
Algorithm 1 LocalUpdate
[…]
1: Set $\mathbf{w}_{1}=\mathbf{w},\alpha_{1}=\alpha,\beta_{1}=\beta$
2:for $p=1,...,P$ do
3:  Sample minibatch $\mathcal{B}_{p}$ randomly from $\mathcal{D}$
4:  Do forward pass with $Q$ on the minibatch $\mathcal{B}_{p}$
5:  Do backwards pass using STE and update $\mathbf{w}_{p+1},\alpha_{p+1},\beta_{p+1}$
6:end for
7: Return $\mathbf{w}_{P+1},\alpha_{P+1},\beta_{P+1}$
@@
where the exponent bias depends on the clipping value $\alpha$ as $b=2^{e}-\log_{2}\alpha+\log_{2}(2-2^{-m})-1$. At training time, $\alpha$ is first initialized using the maximum absolute value of each weight range, and then treated as a learnable parameter that is updated during training. Furthermore, the gradients of the non-differentiable rounding operators are computed using the straight-through estimator (STE) bengio2013estimating , i.e. $\frac{\partial\left\lfloor x_{i}\right\rceil}{\partial x_{i}}=1$, with a key exception being $\log_{2}|x_{i}|$, which is treated as a constant following a similar approach as in kuzmin2022fp8 . Activations are quantized using the same procedure, but with a separate clipping value $\beta$. A summary of the local on-device training procedure is given in Algorithm 1.
Algorithm 2 ServerOptimize
[…]
1: Compute $m_{t}=\sum\limits_{k}n_{k}$
2: Using gradient descent, update the weights as $\mathbf{w}_{t+1}\leftarrow\operatorname*{arg\,min}\limits_{\mathbf{w}}\sum\limits_{k}\frac{n_{k}}{m_{t}}\|Q(\mathbf{w};\alpha_{t+1})-Q(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}$
3: Set $S\leftarrow[\min\limits_{k}\alpha_{t}^{k},\max\limits_{k}\alpha_{t}^{k}]$
4: Using grid search, update the parameters as $\alpha_{t+1}\shortleftarrow\operatorname*{arg\,min}\limits_{\alpha\in S}\sum\limits_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\|Q(\mathbf{w}_{t+1};\alpha)-Q(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}$
5: Return $\mathbf{w}_{t+1},\alpha_{t+1}$
3.4 Unbiased Quantized Communication
@@
When applying FP8 QAT to a federated leaning scenario, an important aspect is the ability to reduce communication overhead by transferring weights between clients and the server using only 8 bits per scalar value. On client devices with hardware for FP8 mixed-precision training support, a copy of high-precision master weights NVMixed is present as in our QAT setup. Therefore, at the end of each communication round, the participating clients need to perform a hard reset of their master weights to the de-quantized FP8 values on a quantization grid. This approach allows for cost reduction in both the uplink and downlink communication.
@@
$Q_{\text{rand}}(x_{i};\alpha)=s_{i}\begin{cases}\left\lceil\frac{x_{i}}{s_{i}}\right\rceil&\kappa\leq\frac{x_{i}}{s_{i}}-\left\lfloor\frac{x_{i}}{s_{i}}\right\rfloor\vskip 1.99997pt\\ \left\lfloor\frac{x_{i}}{s_{i}}\right\rfloor&\text{otherwise,}\end{cases}$ (5)
@@
for $-\alpha\leq x_{i}\leq\alpha$ and where $\kappa$ is a Bernoulli random variable that takes the values 0 and 1 with equal probability. It is straightforward to verify that this randomized quantization is unbiased from a statistics point of view, i.e. for $-\alpha\leq x_{i}\leq\alpha$ we have $\mathbb{E}[Q_{\text{rand}}(x_{i};\alpha)]=x_{i}$, while the deterministic quantization introduced earlier is biased111 The unbiasedness of stochastic quantization assumes a finite domain of the input variable, such that no clipping occurs. While clipping may occur for some values in practice, this does not affect the majority of the weights or activations. In order to simplify theoretical analysis, the assumption of no clipping is often made in the literature, see for example li2017training ..
The quantized weights are then aggregated at the server using a weighted federated average as
$\mathbf{w}_{t+1}\leftarrow\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1}),$ (6)
@@
where $m_{t}=\sum_{{k^{\prime}}\in\mathcal{P}_{t}}n_{k^{\prime}}$. The clipping values are also aggregated, but without quantization, since their contribution to the communication overhead is small relative to the weights. The aggregated weights are then quantized again to FP8 and transmitted to the next set of active clients with a new set of quantization parameters.
An illustrative example of weight quantization is shown in Figure 2a-d, where different weight distributions and the corresponding quantization errors are shown for both deterministic and stochastic quantization. For a given scalar weight, deterministic quantization always results in a lower quantization error, but for communication purposes we are more interested in the quantization error of the aggregated server weights. As can be seen in Figure 2d, stochastic quantization yields a lower error for the average model. For illustration purposes, we have here assumed that the weights are identical on each device, which is not the case in practice. Nevertheless, with stochastic quantization, the average quantized model will be an unbiased estimate of the average unquantized model.
(a) FP32 weight distribution
(b) FP8 weight distribution
(c) Quantization error
distribution.
(d) Quantization error
distribution of average model.
Figure 2: Examples of FP8 quantized weights using the E4M3 format and the corresponding quantization errors for different weight distributions. The range $\alpha$ of the quantization is calculated such that the entire distribution fits within the dynamic range of the FP8 representation. In general, deterministic quantization yields the lowest quantization error for any distribution. However, when weights are quantized independently on several devices (here we use $K=10$) and averaged, stochastic quantization yields smaller errors. Top row: uniform weight distribution, which is a common initialization for convolutional layers. Bottom row: truncated normal distribution, which is commonly used to initialize the linear layers in Transformer models.
3.5 Server-Side Optimization (ServerOptimize)
The standard federated average of the weights in the un-quantized scenario corresponds to the minimization of weighted mean squared error (MSE) between the server parameters and the client parameters. However, when the server parameters are quantized before transmission to the clients, this property no longer holds. We therefore propose a modification to the server-side model aggregation, where the MSE is explicitly minimized. This can be done without communication overhead since all computations are done on the server. Another advantage of this approach is that it leverages the computing power of the server to do more optimization, since the server typically possesses more computing power than a client device.
At time step $t$, we perform model/parameters aggregation to obtain $\mathbf{w}_{t+1},\alpha_{t+1}$ for the next communication round by minimizing the following mean-squared error (MSE).
Note that when there is no communication quantization, the closed-form solution to ServerOptimize is the federated average $\mathbf{w}_{t+1}=\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\mathbf{w}_{t}^{k}$. Since the problem above has no closed-form solution for the quantized communication case, we use the alternating minimization approach to optimize $\mathbf{w}$ and $\alpha$. First, we optimize the model weights $\mathbf{w}$, while fixing $\alpha$ to $\alpha_{t+1}=\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\alpha_{t+1}^{k}$. This is done by minimizing the MSE as
$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}}\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\|Q_{\text{rand}}(\mathbf{w};\alpha_{t+1})-Q_{\text{rand}}(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}.$ (7)
This minimization can be done using gradient descent with a fixed number of iterations
$\alpha_{t+1}=\operatorname*{arg\,min}_{\alpha\in S}\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\|Q_{\text{rand}}(\mathbf{w}_{t+1};\alpha)-Q_{\text{rand}}(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}.$ (8)
The ServerOptimize routine is given in Algorithm 2, which takes place completely on the server and can be used as an optional step in order to improve aggregation of the model weights.
3.6 Overall algorithm
A summary of our proposed FP8 FedAvg with unbiased communication (FP8FedAvg-UQ) method is presented in Algorithm 3, where the optional server-optimization step (UQ+) corresponds to replacing the quantized federated averaging of the weight and range parameters with our two-step MSE minimization optimization in Equations (7) and (8). It is important to note that our method involves two different quantization operators. On-device QAT uses a deterministic and biased quantizer, while the communication part adopts its stochastic counterpart which is unbiased. In the next section, we will give a convergence analysis result for FP8FedAvg-UQ and show that these design choices are well-motivated from a theoretical perspective.